OpenStudy (anonymous):

What is the sum of an 8–term geometric series if the first term is 12 and the last term is –3,359,232?

OpenStudy (lgbasallote):

let A1 = 12 A8 = -3,359,232 n= 8 \[\huge A _{n} = A _{1}r ^{n-1}\] \[\huge -3,359,232 =12r ^{8-1}\] \[\huge -3,359,232 =12r ^{7}\] \[\huge \frac {-3,359,232}{12} =\frac {12r ^{7}}{12}\] \[\huge 279,936=r ^{7}\] \[\huge 6 = r\] now can you do the S_n from here?

OpenStudy (anonymous):

no :(

OpenStudy (lgbasallote):

okay...gimme a few to write it..

OpenStudy (lgbasallote):

do you know the formula for the sum of geometric sequence?

OpenStudy (anonymous):

is it an=a1r(n-1) ?

OpenStudy (lgbasallote):

noo that's for finding a certain term in the series..it's what i used in the above see? the equation is \[\huge S _{n} = \frac {a _{1}(1 - r ^{n})}{1-r}\]

OpenStudy (lgbasallote):

so nowww...we substitute.. A1 = 12 r = 6 n = 8 can you do it now?

OpenStudy (anonymous):

is the answer 4,031,076

OpenStudy (lgbasallote):

yes you're right..sorry bout the confusion :D